# Characterizations of graph classes via convex geometries: A survey

### v6

@@ Abstract
Graph convexity has been used as an important tool to better understand the structure of classes of graphs. Many studies are devoted to determine if a graph equipped with a convexity is a convex geometry. In this work we survey results on characterizations of well-known classes of graphs via convex geometries. We also give some contributions to this subject.
@@ 1 Introduction
A convexity on a nonempty set $V$ is a family $\mathscr{C}$ of subsets of $V$ (called convex sets) such that: (a) $\emptyset$ and $V$ are convex sets; (b) the intersection of convex sets is a convex set. If $V=V(G)$ for some graph $G$ then $\mathscr{C}$ is a graph convexity of $G$. The structure of a convexity is based on the way convex sets are defined. In the context of graph convexities, there are many examples in the literature where convex sets are defined over a path system, typically according to the following rule: fix a family ${\mathscr{P}}_{G}$ of paths (or, more generally, walks) of a graph $G$, and say that a set $S\subseteq V(G)$ is convex if no vertex $x\notin S$ lies in a member of ${\mathscr{P}}_{G}$ that starts and ends at two vertices of $S$. The most natural example is the so-called geodesic convexity pelayo , for which ${\mathscr{P}}_{G}$ is the family of all the shortest paths of $G$. Other important examples are: the monophonic convexity dourado-et-al ; duchet , the $m^{3}$-convexity dragan-et-al , the $l^{k}$-convexity gpt , the toll convexity alcon-et-al , and the weakly toll convexity gutierrez-tondato , defined over induced paths, induced paths of length at least three, induced paths of length at most $k$, tolled walks, and weakly toll walks, respectively. Other rules to define convex sets, not based on path systems, have also been investigated; an important example comes from the so-called Steiner convexity, introduced in caceres . In this case, a set $S\subseteq V(G)$ is Steiner convex if, for every $S^{\prime}\subseteq S$, the vertices of any Steiner tree with terminal set $S^{\prime}$ belong to $S$.
@@
Graph convexities have been studied in many contexts. A strong direction of research has focused on determining convexity invariants, such as the hull number, the interval number, and the convexity number, among others. A major reference work by Pelayo pelayo gives an extensive overview on convexity invariants, applied to the case of the geodesic convexity.
@@
Other studies are devoted to determine if a graph convexity is a convex geometry. This line of research is closely linked to the original idea of defining a combinatorial abstraction of convex sets in Geometry. The convex sets in the plane are those subsets $S$ such that no point outside $S$ lies on a line segment with endpoints in $S$. The convex hull of a set of points $A$ in the plane is the smallest convex set that contains $A$. A point $x$ of a convex set $S$ is an extreme point of $S$ if $S\setminus\{x\}$ is also a convex set, that is, there is no line segment with endpoints in $S\setminus\{x\}$ containing $x$. For instance, an $n$-polygon (considered as a closed region) is clearly a convex set in the plane, and its extreme points are precisely its $n$ vertices [^1] Here, the term vertices is used in the geometric sense.. In addition, the polygon is the convex hull of its $n$ extreme points. Another example is a circle (closed circular region); it is clearly a convex set in the plane, and all the points lying on the circle border are its extreme points. Again, the circle is the convex hull of such points.
@@
The above concepts can be transferred to the combinatorial field in a natural way. We refer the reader to farber-jamison . Let $G$ be a graph and let $\mathscr{C}$ be a convexity of $G$. Given a set $S\subseteq V(G)$, the smallest set $H\in\mathscr{C}$ containing $S$ is called the convex hull of $S$. A vertex $x$ of a convex set $S$ is an extreme vertex of $S$ if $S\backslash\{x\}$ is also convex. The convexity $\mathscr{C}$ is a convex geometry if it satisfies the Minkowski-Krein-Milman property krein-milman : Every convex set is the convex hull of its extreme vertices. The main question dealt with in this survey is: by fixing a rule $r$ to define the convex sets (e.g., a rule based on some path system), determine the class of graphs whose $r$-convexities are convex geometries. For instance, by fixing induced paths, we can obtain the following characterization: a graph $G$ is chordal if and only if the monophonic convexity of $G$ is a convex geometry farber-jamison . Ptolemaic graphs, interval graphs, proper interval graphs, weak polarizable graphs, and 3-fan-free chordal graphs can also be characterized in this way by considering, respectively, the geodesic convexity farber-jamison , the toll covexity alcon-et-al , the weakly toll convexity gutierrez-tondato , the $m^{3}$-convexity dragan-et-al , and the Steiner convexity caceres . In gpt , a characterization of graphs with $l^{k}$-convexities that are convex geometries are studied. Section 3 discusses in detail most of these characterizations.
@@
Convex geometries are equivalent, by complementation, to antimatroids (see korte-lovasz-schrader ). An antimatroid consists of a finite family ${\mathcal{F}}$ of finite sets (called feasible sets) such that: (a) ${\mathcal{F}}$ is closed under unions; (b) if $S$ is a nonempty feasible set, then there is $x\in S$ for which $S\setminus\{x\}$ is also a feasible set (that is, ${\mathcal{F}}$ is an accessible set system). If $U$ is the union of the sets in ${\mathcal{F}}$ then the family of complementary sets ${\mathscr{C}}=\{U\setminus S\mid S\in{\mathcal{F}}\}$ is a convex geometry. It is not difficult to see that ${\mathscr{C}}$ is a convexity; in addition, it satisfies the anti-exchange property edelman ; korte-lovasz-schrader , which is complementary to the accessibility property of antimatroids (property (b) above). In Section 2, we present in more detail the anti-exchange property, which will be useful to derive some results in Section 4.
@@
This survey is organized as follows. Section 2 provides all the necessary background. In Section 3, we review the main results in the literature on characterizations of graph classes via convex geometries; chordal, Ptolemaic, strongly chordal, interval, and proper interval graphs, among other classes, are dealt with. In Section 4 we present some contributions; namely, we show that forests, forests of stars, cographs, bipartite graphs, and planar graphs can be well characterized via convex geometries. Our concluding remarks are presented in Section 5.
@@ 2 Preliminaries
All the graphs in this work are connected, unless otherwise stated. Let ${\mathscr{P}}$ be a function (called path system) that maps each connected graph $G$ to a collection ${\mathscr{P}}_{G}$ of walks of $G$. For instance, ${\mathscr{P}}$ can map $G$ to its shortest paths, induced paths, tolled walks etc. The members of ${\mathscr{P}}_{G}$ are generically called p-walks of $G$. All the path systems dealt with in this paper satisfy the following property: For every $G$, all the walks of $G$ formed by a single vertex or two adjacent vertices are p-walks.
@@
A vertex $x$ of a convex set $S\subseteq V(G)$ is an extreme vertex of $S$ if $S\setminus\{x\}$ is also a convex set. This definition implies that: (a) no p-walk in $G[S]$ contains $x$ as an internal vertex; (b) $x\notin I(u,v)$ for any pair of distinct vertices $u,v\in S$. The set of extreme vertices of $S$ is denoted by ${\mathit{e}xt}(S)$. The convex hull of a set $S\subseteq V(G)$, denoted by $H(S)$, is the smallest set of vertices of $G$ that contains $S$ and is convex. It can be shown that $H(S)$ can be iteratively constructed by successive applications of the interval operator, as follows: starting with a set $S\subseteq V(G)$, define $S_{0}=S$ and $S_{i}=I(S_{i-1}),i\geq 1$. Then there exists $k\geq 1$ such that $S_{k}=S_{k-1}=H(S)$.
@@
We say that a convexity ${\mathscr{C}}$ of a graph $G$ is a convex geometry (or is geometric) if it satisfies the following property: for every $S\in{\mathscr{C}}$, it holds that $S=H({\mathit{e}xt}(S))$. This is called the Minkowski-Krein-Milman property krein-milman . Alternatively, the convexity ${\mathscr{C}}$ is a convex geometry if the following axiom is satisfied:
Anti-exchange property edelman ; korte-lovasz-schrader
If $y,z\notin H(S)$ and $z\in H(S\cup\{y\})$ then $y\notin H(S\cup\{z\})$.
The anti-exchange property is a combinatorial abstraction of the usual convex hull operator in the plane: for two points $y$ and $z$ not in the convex hull of $S$, if $z$ is in the convex hull of $S\cup\{y\}$ then $y$ is outside the convex hull of $S\cup\{z\}$. See Figure 1.
@@ 3 An overview of graph classes and convex geometries
In this section we briefly review the main existing results on convex geometries associated with path systems. Chordal graphs play in important role in this section. For basic facts on chordal graphs, see golumbic80 .
@@ 3.1 Monophonic convexity
A path (cycle) in a graph $G$ is induced if there is no edge of $G$ linking two nonconsecutive vertices of it. Let ${\mathscr{P}}^{m}$ be the path system that maps each graph $G$ to its induced paths. The convexity defined over the induced paths of $G$ is called the monophonic convexity of $G$.
For a vertex $v\in V(G)$, let $N[v]$ denote the closed neighborhood of $v$ in $G$. If $N[v]$ is a clique then $v$ is said to be a simplicial vertex. It is easy to see that $v$ is an extreme vertex of a (monophonically) convex set $S\subseteq V(G)$ if and only if $v$ is a simplicial vertex of $G[S]$, the subgraph of $G$ induced by $S$.
Suppose that $G$ contains an induced cycle $C$ with at least four vertices. Clearly, $S=H(V(C))$ is convex and contains no simplicial vertices, i.e., ${\mathit{e}xt}(S)=\emptyset$. Thus, the Minkowski-Krein-Milman property fails in this case. In other words, if the monophonic convexity of $G$ is a convex geometry then $G$ is chordal. On the other hand, if $G$ is chordal then every nonsimplicial vertex of $G$ lies on an induced path between two simplicial vertices farber-jamison . As chordality is an hereditary property, this implies that every convex subset of vertices of $G$ is indeed the convex hull of its extreme vertices. Hence:
@@
farber-jamison A graph $G$ is chordal if and only if the monophonic convexity of $G$ is a convex geometry.
@@ 3.2 Geodesic convexity
As in the previous subsection, it is easy to see that $v$ is an extreme vertex of a (geodesically) convex set $S\subseteq V(G)$ if and only if $v$ is a simplicial vertex of $G[S]$. In addition, if the geodesic convexity of $G$ is a convex geometry then $G$ must be clearly chordal. However, it is not true that, for every chordal graph $G$, the geodesic convexity of $G$ is a convex geometry. For instance, consider the graph $G^{\prime}$ (called gem) with $V(G^{\prime})=\{a,b,c,d,e\}$ and $E(G^{\prime})=\{ab,bc,cd,de,ae,be,ce,de\}$. Assume that $G^{\prime}$ is an induced subgraph of a chordal graph $G$. Clearly, the extreme vertices of $Y=H(V(G^{\prime}))$ are in $\{a,d\}$. However, $H(\{a,d\})\neq Y$.
A graph $G$ is Ptolemaic if it is chordal and contains no gem as an induced subgraph. The arguments in the previous paragraph tell us that if the geodesic convexity of $G$ is a convex geometry then $G$ is a Ptolemaic graph. Conversely, if $G$ is a Ptolemaic graph then it can be proved that every induced path of $G$ is a shortest path howorka , i.e., the monophonic and the geodesic convexities of $G$ coincide in this case. Therefore, in a Ptolemaic graph, every nonsimplicial vertex of $G$ lies on a shortest path between two simplicial vertices. Consequently, every convex subset of vertices of $G$ is indeed the convex hull of its extreme vertices. Hence:
@@
farber-jamison A graph $G$ is Ptolemaic if and only the geodesic convexity of $G$ is a convex geometry.
@@ 3.3 Strong convexity
An even cycle is a cycle with an even number of vertices. A graph is strongly chordal farber if it is chordal and every of its even cycles with at least six vertices has an odd chord. A vertex of a graph is simple if the neighborhoods of its neighbors form a nested family of sets. Note that a simple vertex is simplicial, but not conversely. In farber , the following characterization of strongly chordal graphs is given:
@@
farber A graph $G$ is a strongly chordal graph if and only if every induced subgraph of $G$ contains a simple vertex.
@@
farber-jamison Let $G$ be a chordal graph. Then a vertex $v\in V(G)$ is an extreme vertex of an strong convex set $S$ if and only if $v$ is a simple vertex in $G[S]$.
By Theorem 3 and Lemma 4, $G$ is a strongly chordal graph. Conversely, suppose that $G$ is a strongly chordal graph. In farber-jamison , the following result is proved:
@@
farber-jamison In a strongly chordal graph, every nonsimple vertex lies in an even-chorded path between simple vertices.
@@
farber-jamison A graph $G$ is strongly chordal if and only if the strong convexity of $G$ is a convex geometry.

### v7

@@ Abstract
Graph convexity has been used as an important tool to better understand the structure of classes of graphs. Many studies are devoted to determine if a graph equipped with a convexity is a convex geometry. In this work, we survey results on characterizations of well-known classes of graphs via convex geometries. We also give some contributions to this subject.
@@ 1 Introduction
A convexity on a nonempty set $V$ is a family $\mathscr{C}$ of subsets of $V$ (called convex sets) such that: (a) $\emptyset$ and $V$ are convex sets; (b) the intersection of convex sets is a convex set. If $V=V(G)$ for some graph $G$, then $\mathscr{C}$ is a graph convexity of $G$. The structure of a convexity is based on the way convex sets are defined. In the context of graph convexities, there are many examples in the literature where convex sets are defined over a path system, typically according to the following rule: fix a family ${\mathscr{P}}_{G}$ of paths (or, more generally, walks) of a graph $G$, and say that a set $S\subseteq V(G)$ is convex if no vertex $x\notin S$ lies in a member of ${\mathscr{P}}_{G}$ that starts and ends at two vertices of $S$. The most natural example is the so-called geodesic convexity [38], for which ${\mathscr{P}}_{G}$ is the family of all the shortest paths of $G$. Other important examples are: the monophonic convexity [17, 20], the $m^{3}$-convexity [19], the $l^{k}$-convexity [27], the toll convexity [2], and the weakly toll convexity [28], defined over induced paths, induced paths of length at least three, induced paths of length at most $k$, tolled walks, and weakly toll walks, respectively. Other rules to define convex sets, not based on path systems, have also been investigated; an important example comes from the so-called Steiner convexity, introduced in [9]. In this case, a set $S\subseteq V(G)$ is Steiner convex if, for every $S^{\prime}\subseteq S$, the vertices of any Steiner tree with terminal set $S^{\prime}$ belong to $S$.
@@
Graph convexities have been studied in many contexts. A strong direction of research has focused on determining convexity invariants, such as the hull number, the interval number, and the convexity number, among others. A major reference work by Pelayo [38] gives an extensive overview on convexity invariants, applied to the case of the geodesic convexity.
@@
Other studies are devoted to determine if a graph convexity is a convex geometry. This line of research is closely linked to the original idea of defining a combinatorial abstraction of convex sets in Geometry. The convex sets in the plane are those subsets $S$ such that no point outside $S$ lies on a line segment with endpoints in $S$. The convex hull of a set of points $A$ in the plane is the smallest convex set that contains $A$. A point $x$ of a convex set $S$ is an extreme point of $S$ if $S\setminus\{x\}$ is also a convex set, that is, if there is no line segment with endpoints in $S\setminus\{x\}$ containing $x$. For instance, an $n$-polygon (considered as a closed region) is clearly a convex set in the plane, and its extreme points are precisely its $n$ vertices [^2] Here, the term vertices is used in the geometric sense.. In addition, the polygon is the convex hull of its $n$ extreme points. Another example is a circle (closed circular region); it is clearly a convex set in the plane, and all the points lying on the circle border are its extreme points. Again, the circle is the convex hull of such points.
@@
The above concepts can be transferred to the combinatorial field in a natural way. We refer the reader to [23]. Let $G$ be a graph and let $\mathscr{C}$ be a convexity of $G$. Given a set $S\subseteq V(G)$, the smallest set $H\in\mathscr{C}$ containing $S$ is called the convex hull of $S$. A vertex $x$ of a convex set $S$ is an extreme vertex of $S$ if $S\backslash\{x\}$ is also convex. The convexity $\mathscr{C}$ is a convex geometry if it satisfies the Krein-Milman (or Minkowski-Krein-Milman) property [34, 37]: Every convex set is the convex hull of its extreme vertices. The main question dealt with in this survey is: by fixing a rule $r$ to define the convex sets (e.g., a rule based on some path system), determine the class of graphs whose $r$-convexities are convex geometries. For instance, by fixing induced paths, we can obtain the following characterization: a graph $G$ is chordal if and only if the monophonic convexity of $G$ is a convex geometry [23]. Ptolemaic graphs, interval graphs, proper interval graphs, weak bipolarizable graphs, and 3-fan-free chordal graphs can also be characterized in this way by considering, respectively, the geodesic convexity [23], the toll covexity [2], the weakly toll convexity [28], the $m^{3}$-convexity [19], and the Steiner convexity [9]. In [27], a characterization of graphs with $l^{k}$-convexities that are convex geometries are studied. Section 3 discusses in detail most of these characterizations.
@@
Convex geometries are equivalent, by complementation, to antimatroids (see [32]). An antimatroid consists of a finite family ${\mathcal{F}}$ of finite sets (called feasible sets) such that: (a) ${\mathcal{F}}$ is closed under unions; (b) if $S$ is a nonempty feasible set, then there is $x\in S$ for which $S\setminus\{x\}$ is also a feasible set (that is, ${\mathcal{F}}$ is an accessible set system). If $U$ is the union of the sets in ${\mathcal{F}}$, then the family of complementary sets ${\mathscr{C}}=\{U\setminus S\mid S\in{\mathcal{F}}\}$ is a convex geometry. It is not difficult to see that ${\mathscr{C}}$ is a convexity; in addition, it satisfies the anti-exchange property [21, 32], which is complementary to the accessibility property of antimatroids (property (b) above). In Section 2, we present in more detail the anti-exchange property, which will be useful to derive some results in Section 4.
@@
This survey is organized as follows. Section 2 provides the necessary background. In Section 3, we review the main results in the literature on characterizations of graph classes via convex geometries; chordal, Ptolemaic, strongly chordal, interval, proper interval graphs, among other classes, are dealt with. In Section 4 we present some contributions; namely, we show that forests, forests of stars, cographs, bipartite graphs, planar graphs, and triangle-free graphs can be well characterized via convex geometries. Our concluding remarks are presented in Section 5.
@@ 2 Preliminaries
All the graphs in this work are connected, unless otherwise stated. A path (resp., cycle) in a graph $G$ is induced if there is no edge of $G$ linking two nonconsecutive vertices of it. We denote by $P_{n}$ (resp., $C_{n}$) the induced path (resp., cycle) with $n$ vertices. For a vertex $v\in V(G)$, let $N(v)$ denote the open neighborhood (or simply neighborhood) of $v$ in $G$. In addition, let $N[v]$ denote the closed neighborhood of $v$ in $G$, where $N[v]=N(v)\cup\{v\}$. If $N[v]$ is a clique, then $v$ is said to be a simplicial vertex. For a set $S\subseteq V(G)$, we denote by $G[S]$ the subgraph of $G$ induced by $S$, defined as follows: $V(G[S])=S$ and $E(G[S])=\{xy\in E(G)\mid x,y\in S\}$. For a graph $F$, we say that $G$ is $F$-free if $G$ does not contain $F$ as an induced subgraph. For a family =${\cal F}=\{G_{1},G_{2},\ldots,G_{k}\}$ of prefixed graphs, we say that $G$ is $(G_{1},G_{2},\ldots,G_{k})$-free (or simply ${\cal F}$-free) if $G$ is $G_{i}$-free for $1\leq i\leq k$. For ${\cal C}=\{C_{k}\mid k\geq 4\}$, $G$ is said to be a chordal graph if $G$ is ${\cal C}$-free.
Let ${\mathscr{P}}$ be a function (called path system) that maps each connected graph $G$ to a collection ${\mathscr{P}}_{G}$ of walks of $G$. For instance, ${\mathscr{P}}$ can map $G$ to its shortest paths, induced paths, tolled walks etc. The members of ${\mathscr{P}}_{G}$ are generically called p-walks of $G$. All the path systems dealt with in this paper satisfy the following property: For every $G$, all the walks of $G$ formed by a single vertex or two adjacent vertices are p-walks.
@@
A vertex $x$ of a convex set $S\subseteq V(G)$ is an extreme vertex of $S$ if $S\setminus\{x\}$ is also a convex set. This definition implies that: (a) no p-walk in $G[S]$ contains $x$ as an internal vertex; (b) $x\notin I(u,v)$ for any pair of distinct vertices $u,v\in S$. The set of extreme vertices of $S$ is denoted by ${\mathit{e}xt}(S)$. The convex hull of a set $S\subseteq V(G)$, denoted by $H(S)$, is the smallest set convex set of $G$ containing $S$. It can be shown that $H(S)$ can be iteratively constructed by successive applications of the interval operator, as follows: starting with a set $S\subseteq V(G)$, define $S_{0}=S$ and $S_{i}=I(S_{i-1}),i\geq 1$. Then, there exists $k\geq 1$ such that $S_{k}=S_{k-1}=H(S)$.
@@
We say that a convexity ${\mathscr{C}}$ of a graph $G$ is a convex geometry (or that it is geometric) if it satisfies the following property: for every $S\in{\mathscr{C}}$, it holds that $S=H({\mathit{e}xt}(S))$. This is called the Krein-Milman property [34]. Alternatively, the convexity ${\mathscr{C}}$ is a convex geometry if the following axiom is satisfied:
Anti-exchange property [21, 32]
If $y,z\notin H(S)$ and $z\in H(S\cup\{y\})$, then $y\notin H(S\cup\{z\})$.
The anti-exchange property is a combinatorial abstraction of the usual convex hull operator in the plane: for two points $y$ and $z$ not in the convex hull of $S$, if $z$ is in the convex hull of $S\cup\{y\}$, then $y$ is outside the convex hull of $S\cup\{z\}$. See Figure 1.
@@ 3 An overview of graph classes and convex geometries
In this section, we briefly review the main existing results on convex geometries in the literature on graph convexity. Chordal graphs play in important role in this section. For basic facts on chordal graphs, see [26].
@@ 3.1 Monophonic convexity
Let ${\mathscr{P}}^{m}$ be the path system that maps each graph $G$ to its induced paths. The convexity defined over the induced paths of $G$ is called the monophonic convexity of $G$. It is easy to see that $v$ is an extreme vertex of a (monophonically) convex set $S\subseteq V(G)$ if and only if $v$ is a simplicial vertex of $G[S]$.
Suppose that $G$ contains an induced cycle $C$ with at least four vertices. Clearly, $S=H(V(C))$ is convex and contains no simplicial vertices, i.e., ${\mathit{e}xt}(S)=\emptyset$. Thus, the Krein-Milman property fails in this case. In other words, if the monophonic convexity of $G$ is a convex geometry, then $G$ is chordal. On the other hand, if $G$ is chordal, then every nonsimplicial vertex of $G$ lies on an induced path between two simplicial vertices [23]. As chordality is an hereditary property, this implies that every convex subset of vertices of $G$ is indeed the convex hull of its extreme vertices. Hence:
@@
[23] A graph $G$ is chordal if and only if the monophonic convexity of $G$ is a convex geometry.
@@ 3.2 Geodesic convexity
As in the previous subsection, it is easy to see that $v$ is an extreme vertex of a (geodesically) convex set $S\subseteq V(G)$ if and only if $v$ is a simplicial vertex of $G[S]$. In addition, if the geodesic convexity of $G$ is a convex geometry, then $G$ must be clearly chordal. However, it is not true that, for every chordal graph $G$, the geodesic convexity of $G$ is a convex geometry. For instance, consider the graph $G^{\prime}$ (called gem) with $V(G^{\prime})=\{a,b,c,d,e\}$ and $E(G^{\prime})=\{ab,bc,cd,de,ae,be,ce,de\}$. See Figure 2. Assume that $G^{\prime}$ is an induced subgraph of a chordal graph $G$. Clearly, the extreme vertices of $Y=H(V(G^{\prime}))$ are in $\{a,d\}$. However, $H(\{a,d\})\neq Y$.
$e$$d$$a$$c$$b$
Figure 2: A gem $G^{\prime}$ with $V(G^{\prime})=\{a,b,c,d,e\}$.
A graph $G$ is Ptolemaic if it is chordal and gem-free. The arguments in the previous paragraph tell us that if the geodesic convexity of $G$ is a convex geometry, then $G$ is a Ptolemaic graph. Conversely, if $G$ is a Ptolemaic graph, then it can be proved that every induced path of $G$ is a shortest path [30], i.e., the monophonic and the geodesic convexities of $G$ coincide in this case. Therefore, in a Ptolemaic graph, every nonsimplicial vertex of $G$ lies on a shortest path between two simplicial vertices. Consequently, every convex subset of vertices of $G$ is indeed the convex hull of its extreme vertices. Hence:
@@
[23] A graph $G$ is Ptolemaic if and only the geodesic convexity of $G$ is a convex geometry.
@@ 3.3 Strong convexity
An even cycle is a cycle with an even number of vertices. A graph is strongly chordal [22] if it is chordal and every of its even cycles with at least six vertices has an odd chord. A vertex of a graph is simple if the neighborhoods of its neighbors form a nested family of sets. Note that a simple vertex is simplicial, but not conversely. In [22], the following characterization of strongly chordal graphs is given:
@@
[22] A graph $G$ is a strongly chordal graph if and only if every induced subgraph of $G$ contains a simple vertex.
@@
[23] Let $G$ be a chordal graph. Then, a vertex $v\in V(G)$ is an extreme vertex of an strong convex set $S$ if and only if $v$ is a simple vertex in $G[S]$.
By Theorem 3 and Lemma 4, $G$ is a strongly chordal graph. Conversely, suppose that $G$ is a strongly chordal graph. In [23], the following result is proved:
@@
[23] In a strongly chordal graph, every nonsimple vertex lies in an even-chorded path between simple vertices.
@@
[23] A graph $G$ is strongly chordal if and only if the strong convexity of $G$ is a convex geometry.